OpenStudy (anonymous):

sqrt7/36 + sqrt252/25

OpenStudy (lgbasallote):

\[\sqrt{\frac{7}{36}} + \sqrt{\frac{252}{25}}?\]

OpenStudy (anonymous):

Yes

OpenStudy (lgbasallote):

\[\implies \frac{\sqrt 7}{\sqrt{36}} +\frac{\sqrt{252}}{\sqrt{25}}\] \[\implies \frac{\sqrt 7}{6} + \frac{\sqrt {252}}{5}\] \[\implies \frac{1}{6} \sqrt 7 + \frac{1}{5} \sqrt{252}\] \[\implies \frac 16 \sqrt 7 + \frac 15 \sqrt{36 \times 7}\] \[\implies \frac{1}{6} \sqrt 7 + \frac{6}{5} \sqrt 7\]

OpenStudy (lgbasallote):

got it?

OpenStudy (anonymous):

Yes Thanks!